OpenStudy (konradzuse):

Write and solve the differential equation that models the verbal statement. (Use k for the constant of proportionality.) The rate of change of y is proportional to y. When x = 0, y = 10, and when x = 2, y = 25. What is the value of y when x = 4?

OpenStudy (agent47):

dy/dt=ky

OpenStudy (konradzuse):

dy/dx = ky y = ?? Evaluate the solution at the specified value of the independent variable. y = ??

OpenStudy (agent47):

wait dy/dx=ky

OpenStudy (agent47):

\[\frac{dy}{y}=kdx\]

OpenStudy (agent47):

Integrate both sides: \[\int{\frac{dy}{y}=\int k*dx}\]

OpenStudy (agent47):

\[ln|y|=kx+C\]

OpenStudy (konradzuse):

I think it's going ot be 25 =10e^k(2)

OpenStudy (agent47):

x = 0, y = 10 ln|10|=C \[ln|y|=kx+ln10\]

OpenStudy (agent47):

\[e^{ln|y|}=e^{kx}*e^{ln10}\] \[y=10e^{kx}\]

OpenStudy (agent47):

x=2, y=25 25=10*e^(2k) 2.5=e^(2k) ln(2.5)=2k k=ln(2.5)/2

OpenStudy (agent47):

\[\large{y=10*e^{\frac{ln2.5}{2}x}}\]

OpenStudy (agent47):

And to figure out what's y when x=4: y=10*e^(2*ln(2.5)), just use a calculator Hopefully I didn't make any silly mistakes

OpenStudy (konradzuse):

nope :).

OpenStudy (agent47):

Yay

OpenStudy (konradzuse):

62.5 yay!

OpenStudy (agent47):

lol when u said nope, I thought the answer didn't wrong... then i saw the smiley and got relieved hahah

OpenStudy (konradzuse):

20 more to go!

OpenStudy (agent47):

OMG

OpenStudy (konradzuse):

jk :P

OpenStudy (agent47):

Btw I have my own hw lol, I wasn't planning on sticking around. Still have to submit this stupid excel project, write a physics lab, and 2 programs. I have no idea why I'm wasting time here.

OpenStudy (konradzuse):

Got a handful :P. These are the toughes! I'm good in Physics, and what kinds of programs? :P

OpenStudy (agent47):

C++

OpenStudy (konradzuse):

oh yeah you were doing the c++ stuffs.

OpenStudy (konradzuse):

still fighting your teacher about the negating spaces? :P

OpenStudy (agent47):

I know it, it's just tedious... and very boring, and no, I gave up on her.. she can't teach at all

OpenStudy (konradzuse):

sux, I learned on my own....

OpenStudy (agent47):

lol I learned with a diff professor, I was retaking the same course basically

OpenStudy (konradzuse):

Lots of bad professors out there... That's why I did it all on my own... You learn much more than the classroom can teach you...